OpenStudy (anonymous):

find the derivative of f(x) = x^4+ x^3- 2x^2-x

OpenStudy (anonymous):

4x^3+3x^2-4x-1

OpenStudy (anonymous):

these are the most basic derivatives... you NEED to be able to do these yourself. do you understand how to do them?

OpenStudy (anonymous):

Yep, I agree Edge.

OpenStudy (anonymous):

ticorn, if your book isn't helping you then try looking at Paul's online math notes. Also, the format for these derivatives is: f(x) = x^n f'(x) = nx^n-1

OpenStudy (anonymous):

ok yeah i got that I was just making sure the answer was correct. Thanks!